# Conformal Lyapunov Optimization: Optimal Resource Allocation under Deterministic Reliability Constraints

### v1

@@ II-F Edge Inference and Queueing Model
All the DUs injected by the EDs into the network are buffered into separate transmission queues. Each $n$-th node keeps a set of $K$ transmission queues, i.e., a different queue for the traffic of each specific ED. Note that an ED can also potentially serve, as an intermediate node, for the traffic of other EDs. Denote as $Q_{n}^{k}(t)$ the state of the queue of the $n$-th node associated to the $k$-th ED, which is measured in number of DUs.
The evolution of the queue is given by
@@
The first term in (20) accounts for the number outgoing DUs, given by $\sum_{(n,m)\in\mathcal{E}}R_{n,m}^{k}(t)$, and for the number, $I_{n}^{s}(t)$, of processed DU if the node is a server, i.e., $n\in\mathcal{S}$. The second term in (20) quantifies the number of task arrivals at EDs $n\in\mathcal{U}$, and the third term corresponds to traffic incoming from other nodes. Since a DU can be processed only if the corresponding queue is not empty, we have the implication
@@
In the setting under study, it is important to keep track not only of the number of DUs in the queues via (20), but also of their identities. To this end, we define the variable $T_{s}^{k}(t)$ as the generation time of the DU at the head of the queue at the server $s$ associated to the ED $k$ at time $t$. In the case of an empty queue we set $T_{s}^{k}(t)=0$ by convention. Figure 2 illustrates the temporal evolutions of the DUs generated at an ED $k$, as well as the corresponding timings of the decisions at a server $s$. Note that the figure considers a simplified situation where the same server $s$ processes all DUs for the $k$-th ED, which is not the most general case.
@@ II-G Performance Metrics
The design goal is to minimize a weighted objective encompassing the transmission energy (17) and the overall imprecision under strict reliability constraints. To this end, we optimize over the sequence of transmission scheduling $\mathbf{R}(t)=\{R_{n,m}^{k}(t)\}_{(n,m)\in\mathcal{E},k\in\mathcal{U}}$, the transmission powers $\mathbf{P}(t)=\{P_{n,m}(t)\}_{(n,m)\in\mathcal{E}}$, and the inference decisions $\mathbf{I}(t)=\{I_{s}^{k}(t)\}_{s\in\mathcal{S},k\in\mathcal{U}}$. As detailed below, we also introduce a sequence of variables $\mathbf{\Theta}(t)=\{\theta^{k}(t)\}_{k=1}^{K}$, one for each ED $k$, which, according to Sec.II-D, are used to define the level of conservativeness applied by the server $s$ when it processes tasks for the $k$-th ED.
@@
$\overline{L}^{k}=\frac{1}{F}\sum_{f=0}^{F-1}\frac{1}{N_{f}^{k}}\sum_{t=fS+1}^{(f+1)S}\sum_{s\in\mathcal{S}}I_{s}^{k}(t)L_{s}^{k}(t)\leq r^{k}+o(1),$ (23)
@@
is the loss accrued by a decision taken at time $t$ by the server $s$ on the task $\tau^{k}(T_{s}^{k}(t))$; the quantity
@@
denotes the number of DUs of the $k$-th ED, whose decisions on have been taken within the $f$-th frame; and the function $o(1)$ tends to zero as $F\to\infty$. Importantly, the constraint defined in (23) must be satisfied deterministically for each run of the optimization protocol. To this end, the network controls the risk tolerance of the decisions made for each ED $k$ via the sequence of variables $\theta^{k}(t)$.
The optimization objective is given by the weighted sum of the transmission energy (17) and of the overall imprecision across all the EDs, i.e.,
@@
where $\eta\geq 0$ represents a multiplier used to explore the energy/precision trade-off. The overall imprecision is given by
@@
denoting the imprecision accrued by the decision taken by the server $s$ on the DU $\tau^{k}(T_{s}^{k}(t))$.
@@ III Conformal Lyapunov Optimization
In this section we describe and analyze the proposed CLO algorithm, which addresses problem (II-H) by integrating LO [8] and O-CRC [12].
@@ III-A An Overview of Conformal Lyapunov Optimization
Figure 3: CLO is organized into frames indexed by $f=0,1,\dots$, each with a fixed duration of $S$ slots. For each time slot $t$ within frame $f$, the transmission powers $\mathbf{P}(t)$ along with the transmission and inference decisions $\mathbf{R}(t)$ and $\mathbf{I}(t)$, respectively, are obtained by following LO with fixed reliability hyperparameters $\mathbf{\Theta}_{f}=\{\theta^{k}_{f}\}_{k=1}^{K}$. The hyperparameters are then updated at the end of each frame so as to address the reliability constraint (23) via O-CRC.
The long-term reliability constraint (II-Ha) cannot be addressed within the classical LO framework. In fact, LO only supports average long-term constraints, and it cannot meet deterministic worst-case long-term constraints of the form (II-Ha). In contrast, O-CRC targets deterministic constraints as in (II-Ha), but it is not designed to tackle optimization problems, focusing instead on inference only.
A key observation is that, if we removed the constraint (II-Ha) from problem (II-H) and we fixed the reliability-controlling variables $\mathbf{\Theta}(t)$, LO would be directly applicable as a solution method to optimize over the remaining variables $\{\mathbf{P}(t),\mathbf{I}(t),\mathbf{R}(t)\}$. Based on this observation CLO tackles the problem (II-H) including the constraint (II-Ha) by applying LO within each frame assuming fixed reliability variables, and then updating the reliability variables at the end of the frame employing a rule inspired by O-CRC.
As depicted in Figure 3. the reliability variables $\{\theta^{k}_{f}=\theta(fS+1)\}_{k=1}^{K}$ are fixed at the beginning of a frame, and LO is applied to address problem (II-H) without the reliability constraint (II-Ha). In order to meet the long-term reliability constraint (II-Ha), the variables $\{\theta^{k}_{f}\}_{k=1}^{K}$ are then updated at the end of the frame by using feedback about the decisions made within the frame. Intuitively, the updates should decrease the variables $\theta^{k}_{f}$ if the decisions for ED $k$ have been too inaccurate in frame $f$, requiring an increase of the level of conservativeness for the inference outputs of ED $k$’s tasks.
In the next subsections, we will detail CLO algorithm, which is detailed also in Algorithm 1. We start by discussing how to update the reliability variables across frames using O-CRC; and then showing how to adapt the LO framework for optimal power allocation, transmission, and inference scheduling within each frame. Finally, we provide a theoretical analysis that proves the effectiveness of the proposed approach.
@@ III-B Updating Reliability Parameters
As overviewed in the previous subsection, the proposed CLO updates the variables $\mathbf{\Theta}_{f}=\{\theta^{k}_{f}\}_{k=1}^{K}$ at the end of frame $f$ to address reliability constraints (II-Ha). CLO assumes the availability of feedback about the average loss accrued by these decisions. This is obtained by summing the losses $L_{s}^{k}(t)$ in (24) for all DUs processed within the slots the $f$-th frame, i.e.,
$\overline{L}_{f}^{k}=\frac{1}{N_{f}^{k}}\sum_{t=fS+1}^{f(S+1)}\sum_{s\in\mathcal{S}}I_{s}^{k}(t)L_{s}^{k}(t).$ (29)
In practice, the feedback (29) may be obtained from the end user or by recording the outcomes of the inference decisions. For instance, for the inference task of predicting the trajectory of an object in motion [44], an observation of the object’s movement can confirm whether the object is included or not in the decision set, yielding the loss $L_{s}^{k}(t)$.
Based on feedback, about all the decisions taken within frame $f$, CLO updates the reliability variables as [12]
$\theta^{k}_{f+1}=\theta^{k}_{f}+\gamma^{k}\mathbbm{1}\{N^{k}_{f}>0\}(r^{k}-\overline{L}^{k}_{f}),$ (30)
where $\gamma^{k}>0$ is the learning rate. By (30), if the reliability constraint $r^{k}$ is violated within the frame $f$, i.e., if $\overline{L}^{k}_{f}>r^{k}$, the variable $\theta^{k}_{f}$ is decreased, i.e., $\theta^{k}_{f+1}\leq\theta^{k}_{f}$. This leads to more conservative, and thus less precise decisions for ED $k$ in the next frame $f+1$. Conversely, when the the reliability constraint is satisfied within the frame $f$, i.e., $\overline{L}^{k}_{f}<r^{k}$, $\theta^{k}_{f}$ is increased by the update (30), prioritizing precision over reliability.
An important remark pertains the impact of the frame size $S$ on the update (30). Indeed, larger frame sizes $S$ entails a more informative feedback (29), since the loss is averaged over a larger number of decisions. On the other hand, larger frames increase the overall number of time slots required before the updates (30) converge to a stable solution, satisfying the reliability constraint (II-Ha).
The resulting tension between informativeness of each update and number of updates performed within a given number of slots will be studied theoretically in sec. IV.
@@ III-C Within-Frame Optimization of Power Allocation and Transmission/ Inference Scheduling
Specifically, at each time $t$, LO addresses the instantaneous problem (Derivations are detailed in Section I of the supplemental materials)
@@
where $V>0$ is a hyperparameter; and
@@
The objective function in (31) is a weighted sum of the current contribution $F_{\mathrm{tot}}(t)$ to the objective function in the original problem (II-H) and of two penalty terms. The first term, $\sum_{(n,m)\in\mathcal{E},k\in\mathcal{U}}U_{n,m}^{k}(t)R_{n,m}^{k}(t)$, favors transmission for traffic with largest differential backlog [45]. In a similar way, the second term $\sum_{n\in\mathcal{N},k\in\mathcal{U}}\mathbbm{1}\{n\in\mathcal{S}\}Q_{n}^{k}(t)I_{n}^{k}(t)$, favors processing for the servers with the largest number of queued DUs.
@@
III-D Modeling the Imprecision Function
Figure 4: Illustration of the training process for the predictor of the imprecision function associated with the classifier deployed at the server $s$ for an image classification task. The imprecision predictor is given as input the pair $(x,\theta)$ of input $x$ and threshold $\theta$, producing an estimate of the imprecision function value $\hat{F}_{s}(x,\theta)$ associated to the classifier decision. The training loss $\mathcal{L}(F_{s},\hat{F}_{s})$ evaluates the mismatch between the actual and the predicted imprecision.
By their definitions (8), (3) and (5), both the reliability and the imprecision criteria associated with an inference task $\tau$ can be evaluated only after the execution of the task. This is not an issue for the reliability loss function $L_{s}(\tau,\theta)$. In fact, the O-CRC update (30) only requires feedback after a decision is implemented. In contrast, the precision function is requested to solve the instantaneous problem (31), which has to provide the decision variables $I_{s}^{k}(t)$ over the time slots $t$. This requires an estimate of the imprecision function prior to processing the inference task.
To tackle this issue, as illustrated in Figure 4, we propose to train $|\mathcal{S}|$ neural networks devoted to predict the imprecision associated to a pair $(\tau,\theta)$ for each of the $|\mathcal{S}|$ servers. Specifically, the $s$-th neural network corresponds to the inference model employed by the $s$-th server. The trained imprecision predictors $\{F_{s}(\tau,\theta)\}_{s=1}^{|\mathcal{S}|}$ act as approximations of the imprecision functions (28), and they can be employed to evaluate the cost function of the instantaneous optimization problem (31). As depicted in Figure 4, a possible approach consists to train the imprecision predictor on an augmented training set of the original inference task, where we consider a set of possible values for the reliability variable $\theta$ for each training sample. The output variable is represented by the imprecision accrued by each training pair by the $s$-th inference model.
An alternative approach involves training a set of low-complexity networks through knowledge-distillation techniques [46]. In this setup, inference models serve as teacher networks, while precision loss approximators act as student networks. The student models are trained to mimic the outputs of the inference models, thus allowing to obtain a reliable estimate the effective loss. For example, in the context of prediction set construction for image classification, a practical measure of imprecision can be obtained by counting the number of classes for which the student model assigns a confidence level exceeding a predefined threshold.

### v5

@@ II-F Edge Inference and Queueing Model
We assume that the DUs injected by the EDs into the network are buffered into separate transmission queues. Specifically, the $n$-th node has a dedicated queue $Q_{n}^{k}(t)$ for the traffic of the $k$-th ED, which reflects the number of queued DUs. Note that an ED can also potentially serve, as an intermediate node, for the traffic of other EDs.
The evolution of each queue is given by
@@
For each time slot $t$, the queue is updated by subtracting the number of outgoing DUs, given by $\sum_{(n,m)\in\mathcal{E}}R_{n,m}^{k}(t)$, and, if node $n$ is a server (i.e., $n\in\mathcal{S}$), by the number of processed DUs, $I_{n}^{s}(t)$. Conversely, it is incremented by the number of task arrivals at the ED, if $n\in\mathcal{U}$, and by the DUs received from other nodes. Since a DU can be processed only if the corresponding queue is not empty, we have the implication
@@
In the setting under study, it is important to keep track not only of the number of DUs in the queues via (20), but also of their identities. To this end, we define the variable $T_{s}^{k}(t)$ as the generation time of the DU at the head of the queue of the $s$-th server, associated to the $k$-th ED, at time $t$. When the queue is empty we simply set $T_{s}^{k}(t)=0$. Figure 2 illustrates the temporal evolutions of the DUs possibly generated at the $k$-th ED, as well as the corresponding timings of the decisions at the $s$-th server. Note that, for simplicity, the figure considers a simplified situation in which all the DUs of the $k$-th ED are processed by the same server $s$, which is not the general case.
@@ II-G Performance Metrics
The design goal is to minimize a weighted objective encompassing the transmission energy (17) and the overall precision loss, under strict reliability constraints. To this end, we optimize over the sequence of transmission scheduling $\mathbf{R}(t)=\{R_{n,m}^{k}(t)\}_{(n,m)\in\mathcal{E},k\in\mathcal{U}}$, the transmission powers $\mathbf{P}(t)=\{P_{n,m}(t)\}_{(n,m)\in\mathcal{E}}$, and the task assignments $\mathbf{I}(t)=\{I_{s}^{k}(t)\}_{s\in\mathcal{S},k\in\mathcal{U}}$. As detailed below, we also introduce a sequence of variables $\mathbf{\Theta}(t)=\{\theta^{k}(t)\}_{k=1}^{K}$, one for each ED, that, according to Section II-D, are used to define the level of conservativeness applied by the server $s$ when it processes tasks for the $k$-th ED.
@@
$\overline{L}^{k}=\frac{1}{F}\sum_{f=0}^{F-1}\frac{1}{N_{f}^{k}}\sum_{t=fS+1}^{(f+1)S}\sum_{s\in\mathcal{S}}I_{s}^{k}(t)L_{s}^{k}(t)\leq r^{k}+\mathcal{O}\left(\frac{1}{F}\right),$ (23)
@@
is the loss accrued by a decision taken at time $t$ by the server $s$ on the task $\tau^{k}(T_{s}^{k}(t))$; the function $\mathcal{O}(\frac{1}{F})$ tends to zero as $F\to\infty$; and the quantity
@@
denotes the number of DUs of the $k$-th ED, whose decisions on have been taken within the $f$-th frame. Importantly, the constraint defined in (23) must be satisfied deterministically for each run of the optimization protocol. To this end, the network controls the risk tolerance of the decisions made for each ED $k$ via the sequence of variables $\theta^{k}(t)$.
The optimization objective is given by the weighted sum of the transmission energy (17) and of the overall precision loss across all the EDs, i.e.,
@@
where $\eta\geq 0$ is a multiplier used to explore the energy/precision trade-off. The overall precision loss is given by
@@
denoting the precision loss accrued by the decision taken by the server $s$ on the DU $\tau^{k}(T_{s}^{k}(t))$.
@@ III Conformal Lyapunov Optimization
In this section, we describe and analyze the proposed CLO algorithm, which addresses problem (II-H) by integrating LO [8] and O-CRC [12].
@@ III-A An Overview of Conformal Lyapunov Optimization
Figure 3: CLO frame-based structure: each frame $f\in\mathbb{N}_{0}$ is composed by $S$ slots, with fixed duration. In each time slot $t$ within the $f$-th frame, the powers $\mathbf{P}(t)$, the scheduling $\mathbf{R}(t)$, and server assignments $\mathbf{I}(t)$ are obtained by LO for fixed reliability parameters $\mathbf{\Theta}_{f}=\{\theta^{k}_{f}\}_{k=1}^{K}$, which are updated at the end of each frame by O-CRC, to address the reliability constraint (23) .
Classical LO only supports statistically-average long-term constraints, while it cannot address deterministic (worst-case) long-term reliability constraints of the form (II-Ha). Conversely, O-CRC targets deterministic constraints as in (II-Ha), but it is not designed to tackle optimization problems, focusing instead only on inference reliability. The key difference between a statistical average constraint and a deterministic reliability constraint is that the former is satisfied on average across multiple runs of the optimization procedure. Therefore, it is generally violated in any specific run of the system. In contrast, a deterministic reliability constraint is more stringent, as it demands that the given reliability condition be satisfied for each individual run of the procedure.
A key observation is that, if we removed the constraint (II-Ha) from problem (II-H) and we fixed the reliability-controlling variables $\mathbf{\Theta}(t)$, LO would be directly applicable as a solution method to optimize over the remaining variables $\{\mathbf{P}(t),\mathbf{I}(t),\mathbf{R}(t)\}$. Based on this observation, CLO tackles the problem (II-H) including the constraint (II-Ha) by applying LO within each frame assuming fixed reliability variables, and then updating the reliability variables at the end of the frame employing a rule inspired by O-CRC.
As shown in Figure 3, the reliability variables $\{\theta^{k}_{f}=\theta(fS+1)\}_{k=1}^{K}$ are fixed at the beginning of a frame, and LO is applied to address problem (II-H) without the reliability constraint (II-Ha). In order to meet the long-term reliability constraint (II-Ha), the variables $\{\theta^{k}_{f}\}_{k=1}^{K}$ are then updated at the end of the frame by using feedback about the decisions made within the frame. Intuitively, the updates should decrease the variables $\theta^{k}_{f}$ if the decisions for the $k$-th ED have been too inaccurate during the $f$-th frame, requiring an increase of the conservativeness for the inference outputs of the $k$-th ED’s tasks.
The next subsections will provide deeper insights on the CLO algorithm, which is also detailed in Algorithm 1. We start outlining how to update the reliability variables across frames according to O-CRC; and then showing how to adapt the LO framework for optimal power, transmission scheduling, and inference allocations, within each frame. Finally, we provide a theoretical analysis that proves the effectiveness of the proposed approach.
@@ III-B Updating Reliability Parameters
As overviewed in the previous subsection, the proposed CLO updates the variables $\mathbf{\Theta}_{f}=\{\theta^{k}_{f}\}_{k=1}^{K}$ at the end of each frame $f\in\mathbb{N}_{0}$ to address reliability constraints (II-Ha). CLO assumes the availability of feedback about the average loss accrued by these decisions with a delay of $d^{k}$ time-slots for each device $k$. This delay may result from the overhead associated with estimating and disseminating frame loss information. Accordingly, the update of reliability hyperparameters at frame $f$ is based on the average reliability loss observed at frame $f-d^{k}$ for each device $k$.
The average loss at frame $f$ is obtained by summing the losses $L_{s}^{k}(t)$ in (24) for all DUs processed within the slots of the $f$-th frame, i.e.,
$\overline{L}_{f}^{k}=\frac{1}{N_{f}^{k}}\sum_{t=fS+1}^{(f+1)S}\sum_{s\in\mathcal{S}}I_{s}^{k}(t)L_{s}^{k}(t).$ (29)
In practice, the feedback (29) may be obtained by recording the outcomes of the inference decisions. For instance, for the inference task of predicting the trajectory of an object in motion, the subsequent observation of the object’s movement can confirm whether the object pixels are included or not in the decision set, yielding the loss $L_{s}^{k}(t)$ [44].
Based on the received feedback at frame $f$, CLO updates the reliability variables as [12]
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\theta^{k}_{f+1}=\theta^{k}_{f}+\gamma^{k}\mathbbm{1}\{N^{k}_{f-d^{k}}>0\}(r^{k}-\overline{L}^{k}_{f-d^{k}}),}$ (30)
where $\gamma^{k}>0$ is the learning rate. By (30), if the reliability constraint $r^{k}$ is violated within the $(f-d^{k})$-th frame, i.e., if $\overline{L}^{k}_{f-d^{k}}>r^{k}$, the variable $\theta^{k}_{f}$ is decreased, i.e., $\theta^{k}_{f+1}\leq\theta^{k}_{f}$. This leads to more conservative, and thus less precise, decisions for the $k$-th ED during the next $(f+1)$-th frame. Conversely, when the reliability constraint is satisfied within the $(f-d^{k})$-th frame, i.e., $\overline{L}^{k}_{f-d^{k}}<r^{k}$, the parameter $\theta^{k}_{f}$ is increased by the update (30), prioritizing precision over reliability.
An important remark pertains the impact of the frame size $S$ on the update (30). Indeed, larger frame sizes $S$ entails a more informative feedback (29), since the loss is averaged over a larger number of decisions. On the other hand, having larger frames, thus a less frequent update of $\theta^{k}_{f}$, will proportionally increase the overall number of time slots before the updates (30) will converge to a stable solution, satisfying the reliability constraint (II-Ha).
The resulting tension between informativeness of each update and update rate (i.e., convergence rate) will be studied theoretically in Section IV.
@@ III-C Within-Frame Optimization of Power Allocation and Transmission/ Inference Scheduling
Specifically, at each time $t$, LO addresses the instantaneous problem  [^2] Derivations are detailed in Section I of the supplemental material.
@@
where $V>0$ is a hyperparameter that trades energy consumption and precision, for queues congestion (average delay), and
@@
The objective function in (31) is a weighted sum of the current contribution $F_{\mathrm{tot}}(t)$ to the objective function in the original problem (II-H), and two penalty terms. The first term $\sum_{(n,m)\in\mathcal{E},k\in\mathcal{U}}U_{n,m}^{k}(t)R_{n,m}^{k}(t)$, favors transmission for traffic with largest differential backlog [45]. The second term $\sum_{n\in\mathcal{N},k\in\mathcal{U}}\mathbbm{1}\{n\in\mathcal{S}\}Q_{n}^{k}(t)I_{n}^{k}(t)$ favors processing for the servers with the largest number of queued DUs.
@@
III-D Modeling the Precision Loss
Figure 4: Training of the NN predictor of the precision loss function ${F}_{s}(x,\theta)$ associated with the classifier employed at the $s$-th server, for an image classification task. The classifier input $x$ and threshold $\theta$ are the input pair $(x,\theta)$ for the predictor, which produces an estimate $\hat{F}_{s}(x,\theta)$ of the precision loss value possibly associated to the classifier decision. The training loss $\mathcal{L}(F_{s},\hat{F}_{s})$ evaluates the mismatch between the actual and the predicted precision loss.
By their definitions in (8), (3) and (5), both the reliability and the precision losses associated with an inference task $\tau$, can be evaluated only after the execution of the task. This is not an issue for the reliability loss function $L_{s}(\tau,\theta)$. In fact, the O-CRC update (30) only requires feedback after a decision is implemented. In contrast, the precision loss is requested to solve the instantaneous problem (31), which has to provide the decision variables $I_{s}^{k}(t)$ over the time slots $t$. In practice, this requires an estimate of the precision loss function before processing the inference task.
To tackle this issue, as illustrated in Figure 4, we propose to train $|\mathcal{S}|$ neural networks (NNs) devoted to predict the precision loss associated to a pair $(\tau,\theta)$ for each of the $|\mathcal{S}|$ servers. Specifically, the $s$-th NN predictor is associated with the inference model employed by the $s$-th server. The trained predictors $\{\hat{F}_{s}(\tau,\theta)\}_{s=1}^{|\mathcal{S}|}$ act as approximators of the actual precision losses (28), and they can be employed to evaluate the cost function of the instantaneous optimization problem (31). As depicted in Figure 4, a possible approach consists in training the precision loss predictor on an augmented training set of the original inference task, where we consider a set of possible values for the reliability variable $\theta$ for each training sample $\tau$. The output variable is represented by the precision loss accrued by each training pair $(\tau,\theta)$ by the actual $s$-th inference model.
An alternative approach involves training a set of low-complexity networks through knowledge-distillation techniques [46]. In this setup, the actual inference models at the servers, play the role of teacher networks, while precision loss approximators act as student networks. The student models are trained to mimic the outputs of the inference models, thus allowing to obtain a reliable estimate of the effective loss. For example, in the context of prediction-set construction for image classification, a practical measure of imprecision can be obtained by counting the number of classes for which the student model assigns a confidence level exceeding a predefined threshold (see Figure 4).
We note that alternative approximation techniques can also be considered, each with a distinct impact on the algorithm’s performance. The effect of neural network based precision loss approximation is evaluated in Appendix D.